# FAILURE OF LICHNEROWICZ-TYPE RESULT IN PARABOLIC GEOMETRIES OF REAL RANK AT LEAST 33

### v3

@@ 1 Introduction
[…]
@@
Riemannian conformal manifolds are examples of regular, normal, parabolic Cartan geometries. They are modeled on the conformal sphere, a homogeneous space for the action of the orthogonal group $\mathrm{O}(n)$ by Möbius transformations, whose stabilizer is a parabolic subgroup.
CR manifolds satisfy a similar Lichnerowicz-type result. A pseudoconvex CR manifold is a $2n+1$ dimensional manifold $M$ with a $2n$ dimensional co-oriented distribution $H\subset TM$ admitting a complex structure $J:H\to H$ and satisfying certain positive definiteness and integrability conditions. The subbundle $H$ induces an equivalence class of contact forms $\theta]$ vanishing on $H.$ These satisfy the property that $d\theta:H\times H\to\mathbb{R}$ is the imaginary part of a positive definite Hermitian form. Call a CR manifold $(M,H,J)$ essential if $\mathrm{Aut}(M,\theta,J)\subsetneq\mathrm{Aut}(M,H,J)$ for all contact forms $\theta$ in the equivalence class. Then results of Webster [19] and Schoen [18] imply the following.
@@
To make the connection to Cartan geometry, we must generalize the CR condition slightly to partially integrable, almost CR. A pseudoconvex, partially integrable, almost CR manifold is equivalent to a choice of regular, normal, parabolic Cartan geometry modeled on the CR sphere $S^{2n+1}$, a homogeneous space for the action of $\mathrm{SU}(n+1,1)$ by CR diffeomorphisms. In particular, the group $\mathrm{SU}(n+1,1)$ has real rank one and its action on the sphere has a parabolic stabilizer subgroup. Frances [7] generalized the Ferrand-Obata and Schoen-Webster theorems, proving a Lichnerowicz-type result that applies to all real rank one parabolic geometries. Alt [1] modified Frances’s result to prove the following.
@@
There are four series of real rank one parabolic geometries, corresponding to (1) conformal, (2) pseudoconvex, partially integrable, almost CR, (3) quaternionic contact, and (4) octonionic contact structures. These geometric structures are modeled on the homogeneous spaces (1) $\partial\mathbf{H}^{n+1}_{\mathbb{R}},$ (2) $\partial\mathbf{H}^{n+1}_{\mathbb{C}}$, (3) $\partial\mathbf{H}^{n+1}_{\mathbb{H}}$, and (4) $\partial\mathbf{H}^{2}_{\mathbb{O}}$, the boundaries of real, complex, quaternionic, and octonionic hyperbolic spaces.
@@
Lorentzian conformal manifolds of dimension $n\geq 3$ are regular, normal, parabolic Cartan geometries modeled on a homogeneous space for the group $\mathrm{O}(n,2).$ Since $\mathrm{O}(n,2)$ for $n\geq 3$ is simple of real rank $2,$ Lorentzian Lichnerowicz is an intermediate case between Frances and Alt’s Theorem 1.3 for real rank $1$ - a domain where Lichnerowicz-type results hold - and our Theorem 1.4 for real rank at least $3$ - a domain where they fail.
@@ Acknowledgments
2 Background
2.1 Notation
Given a real or complex vector space $V$, we denote its complexification and realification denoted by $V^{\mathbb{C}}$ and $V^{\mathbb{R}}$, respectively. Given an inner product, a vector $v$ determines a covector $v_{\flat}$ and a covector $\alpha$ determines a vector $\alpha^{\sharp}$ via the musical isomorphisms. If $\alpha$ is a root of a Lie algebra $\mathfrak{g}$, $\eta_{\alpha}$ refers to some nonzero vector in the rootspace $\mathfrak{g}_{\alpha}.$ The notation $\alpha_{k}$ refers to a simple root, while $\beta_{k}$ refers to a simple restricted root. For simple roots $\alpha_{i},\alpha_{j},$ the notation $(ij)$ refers to the composition of simple root reflections $s_{i}\circ s_{j}.$ The letter $\mu$ always refers to the highest root of a Lie algebra. Coefficients of the Cartan matrix are written $c_{ij}.$
2.2 Structure Theory
@@
$\mathfrak{g}=Z(\mathfrak{a})\oplus\bigoplus_{\alpha\in\hat{\Delta}}\mathfrak{g}_{\alpha}$
@@
Given a simple system $\Delta^{0}$ for $\Delta^{+}$, the restrictions to $\mathfrak{a}$ of the roots in $\Delta^{0}\backslash\Delta_{c}$ form a simple system $\hat{\Delta}^{0}$ for $\hat{\Delta}^{+},$ with some pairs of simple roots in $\Delta^{0}$ restricting to a single simple restricted root in $\hat{\Delta}^{0},$ and compact simple roots restricting to $0.$ A subset of simple restricted roots $\hat{I}\subset\hat{\Delta}^{0}$ determines a parabolic subalgebra by the following process. Given a restricted root
@@
$\mathfrak{p}=Z(\mathfrak{a})\oplus\bigoplus_{\alpha\in\hat{\Delta}:h_{\hat{I}}(\alpha)\geq 0}\mathfrak{g}_{\alpha}.$
Up to inner automorphism, this gives every parabolic subalgebra of $\mathfrak{g}.$ Therefore we will assume without loss of generality that a given parabolic subalgebra is in this form. Define
@@
$\mathfrak{g}_{0}:=Z(\mathfrak{a})\oplus\bigoplus_{\alpha\in\hat{\Delta}(\mathfrak{g}_{0})}\mathfrak{g}_{\alpha},$
isomorphic to the quotient $\mathfrak{p}/\mathfrak{p}_{+}$ of $\mathfrak{p}$ by its nilradical. Since $\mathfrak{g}_{0}$ is reductive, its derived subalgebra $\mathfrak{g}_{0}^{ss}:=[\mathfrak{g}_{0},\mathfrak{g}_{0}]$ is semisimple.
@@
It follows from Proposition 2.1 that $\beta_{k}^{\sharp}\in\mathfrak{g}_{0}^{ss}\cap\mathfrak{a}.$ The parabolic $\mathfrak{p}=\mathfrak{g}_{0}\oplus\mathfrak{p}^{+}.$ In fact, $h_{\hat{I}}$ induces a grading $\mathfrak{g}=\mathfrak{g}_{-k}\oplus\cdots\oplus\mathfrak{g}_{0}\oplus\cdots\oplus\mathfrak{g}_{k}$, and we define $\mathfrak{g}_{-}:=\mathfrak{g}_{-k}\oplus\cdots\oplus\mathfrak{g}_{-1}$ and $\mathfrak{g}_{+}:=\mathfrak{p}_{+}=\mathfrak{g}_{1}\oplus\cdots\oplus\mathfrak{g}_{k}.$ There exists a special element $E\in\mathfrak{a}$ such that $[E,X]=iX$ for $X\in\mathfrak{g}_{i},$ known as the grading element. In particular, this definition implies $E\in\mathfrak{z}(\mathfrak{g}_{0}).$
@@
for integers $n_{\beta}$. Then $\alpha|_{\mathfrak{a}}=\overline{\alpha}|_{\mathfrak{a}}.$ To draw the Satake diagram for a Lie algebra $\mathfrak{g},$ take the Dynkin diagram for $\mathfrak{g}^{\mathbb{C}},$ color the elements of $\Delta_{c}\cap\Delta^{0}$ black, and for noncompact simple roots $\alpha,$ if $\alpha\neq\overline{\alpha},$ connect $\alpha$ and $\overline{\alpha}$ by a bi-directional arrow. The resulting diagram is independent of a choice of maximally noncompact Cartan subalgebra $\mathfrak{c}$ and positive system $\Delta^{+}\subset\Delta$ compatible with $\sigma$. A full list of Satake diagrams of real Lie algebras $\mathfrak{g}$ such that $\mathfrak{g}^{\mathbb{C}}$ is simple is given in Appendix A.
@@ 2.3 Real and Complex Representations
If $\mathfrak{g}$ is complex semisimple and $X_{1},\dots X_{n}$ is a $\mathbb{C}$-basis for $\mathfrak{g},$ then $X_{1},\dots X_{n},iX_{1},\dots iX_{n}$ is an $\mathbb{R}$-basis for $\mathfrak{g}^{\mathbb{R}}.$ Then given a $\mathbb{C}$-linear endomorphism $T:\mathfrak{g}\to\mathfrak{g},$ we have $\mathrm{tr}_{\mathbb{R}}(T)=2\mathrm{Re}(\mathrm{tr}_{\mathbb{C}}(T)).$ If $\mathfrak{g}$ is a complex Lie algebra, let $\langle\cdot,\cdot\rangle_{\mathbb{C}}$ be the Killing bracket on $\mathfrak{g}$, and let $\langle\cdot,\cdot\rangle_{\mathbb{R}}$ be the Killing bracket on $\mathfrak{g}^{\mathbb{R}}$. Then $\langle x,y\rangle_{\mathbb{R}}=2\mathrm{Re}(\langle x,y\rangle_{\mathbb{C}}).$ Let $\mathfrak{h}_{0}\leq\mathfrak{h}$ be the subspace on which roots of $\mathfrak{g}$ are real-valued. Because $\langle\cdot,\cdot\rangle_{\mathbb{C}}$ is real-valued on $\mathfrak{h}_{0}$, if $x,y\in\mathfrak{h}_{0},$ then $\langle x,y\rangle_{\mathbb{R}}=2\langle x,y\rangle_{\mathbb{C}}.$ These brackets induce isomorphisms $\flat_{\mathbb{C}}:\mathfrak{h}_{0}\mapsto\mathfrak{h}_{0}^{*}$ and $\flat_{\mathbb{R}}:\mathfrak{h}_{0}\to\mathfrak{h}_{0}^{*},$ which both induce Killing brackets on $\mathfrak{h}_{0}^{*}.$ We know $\flat_{\mathbb{R}}=2\flat_{\mathbb{C}}.$ Calling their inverses $\sharp_{\mathbb{R}}$ and $\sharp_{\mathbb{C}},$ $\sharp_{\mathbb{C}}=2\sharp_{\mathbb{R}}.$
@@
Let $,\langle\cdot,\cdot\rangle_{\mathbb{C}}$$\langle\cdot,\cdot\rangle_{\mathbb{R}}$ be the brackets induced on $\mathfrak{h}_{0}^{*}$ by $\mathfrak{g}$ and $\mathfrak{g}^{\mathbb{R}}$ respectively. Then $\langle x,y\rangle_{\mathbb{C}}=2\langle x,y\rangle_{\mathbb{R}}.$
@@
Given $\mathfrak{g}$ real semisimple and a representation $\mathfrak{g}\curvearrowright V$, a vector $v\in V$ is called a weight vector if there exists $\alpha\in\mathfrak{a}^{*}$ such that $H\cdot v=\alpha(H)v$ for all $H\in\mathfrak{a}.$ The weight space $V_{\alpha}$ is the set of such vectors. A weight vector $v$ is called a lowest weight vector if it is annihilated by all negative restricted rootspaces of $\mathfrak{g}.$
Suppose $\mathfrak{g}$ is complex semisimple. Then $\mathfrak{g}^{\mathbb{R}}$ is real semisimple. There exists a compact real form $\mathfrak{u}\leq\mathfrak{g}$ for which the conjugation $\theta_{\mathfrak{u}}$ of $\mathfrak{g}$ about $\mathfrak{u}$ preserves $\mathfrak{h}.$ Then $\theta_{\mathfrak{u}}$ is a Cartan involution of $\mathfrak{g}^{\mathbb{R}}$, and $\mathfrak{h}\leq\mathfrak{g}^{\mathbb{R}}$ is a $\theta_{\mathfrak{u}}$-stable Cartan subalgebra. Any Cartan subalgebra $\mathfrak{c}\leq\mathfrak{g}^{\mathbb{R}}$ is also a Cartan subalgebra of $\mathfrak{g},$ and thus is unique up to inner automorphism. It follows that $\mathfrak{h}\leq\mathfrak{g}^{\mathbb{R}}$ is automatically maximally noncompact. The noncompact part of $\mathfrak{h}$ is $\mathfrak{h}\cap i\mathfrak{u}=\mathfrak{h}_{0},$ the subspace of $\mathfrak{h}$ on which all roots of $\mathfrak{g}$ are real-valued. Then the restricted roots of $\mathfrak{g}^{\mathbb{R}}$ are exactly the roots of $\mathfrak{g}$ restricted to $\mathfrak{h}_{0},$ and each restricted root space $\mathfrak{g}_{\alpha}$ for $\alpha\in\hat{\Delta}$ is a $2$ dimensional real space. Furthermore, if $\mathfrak{g}\curvearrowright V$ is a representation, the restricted weights of $\mathfrak{g}^{\mathbb{R}}\curvearrowright V$ are the weights of $\mathfrak{g}\curvearrowright V$ restricted to $\mathfrak{h}_{0}.$
Given $\mathfrak{g}$ real semisimple, define the fundamental weights $\lambda_{i}\in\mathfrak{h}^{*}$ of $\mathfrak{g}^{\mathbb{C}}$ by the property $\langle\lambda_{i},\alpha_{j}\rangle=\delta_{ij}\frac{|\alpha_{j}|^{2}}{2}$ for all simple roots $\alpha_{j}\in\Delta^{0}.$ Then given an integral weight $\gamma\in\mathfrak{h}^{*}$, we can decompose it as $\gamma=\sum\gamma^{i}\lambda_{i}$ for integers $\gamma^{i}.$ Furthermore, we define restricted fundamental weights $\hat{\lambda}_{i}\in\mathfrak{a}^{*}$ by the property $\langle\hat{\lambda}_{i},\beta_{j}\rangle=\delta_{ij}\frac{|\beta_{j}|^{2}}{2}$ for all simple restricted roots $\beta_{j}\in\hat{\Delta}^{0}$. Given a restricted weight $\sigma,$ we can decompose it as $\sigma=\sum\sigma^{i}\hat{\lambda}_{i}.$

### v4

@@ 1 Introduction
A conformal manifold $(M,c)$ is called essential if, for any metric $g$ in the conformal class $c$, $\mathrm{Isom}(g)\subsetneq\mathrm{Conf}(c)$. That is, the full conformal automorphism group is larger than the isometry group of any representative metric. As conjectured by Lichnerowicz and proven independently by Lelong-Ferrand [10] and Obata [14], there are remarkably few essential conformal manifolds.
@@
Riemannian conformal manifolds are the regular, normal, parabolic Cartan geometries modeled on the conformal sphere. This is the homogeneous space for the action of the orthogonal group $\mathrm{O}(n)$ by Möbius transformations. This model geometry has stabilizer a parabolic subgroup of $O(n).$
CR manifolds satisfy an analogous result to Ferrand-Obata. We will call such results Lichnerowicz-type. A pseudoconvex CR manifold is a $2n+1$ dimensional manifold $M$ with a $2n$ dimensional co-oriented distribution $H\subset TM$ admitting an almost complex structure $J:H\to H$ for which $J^{2}=-1$ and satisfying certain positive definiteness and integrability conditions. The subbundle $H$ induces an equivalence class of contact forms $\theta]$ vanishing on $H.$ These satisfy the property that $d\theta:H\times H\to\mathbb{R}$ is the imaginary part of a positive definite Hermitian form. Call a CR manifold $(M,H,J)$ essential if $\mathrm{Aut}(M,\theta,J)\subsetneq\mathrm{Aut}(M,H,J)$ for all contact forms $\theta$ in the equivalence class. Then results of Webster [19] and Schoen [18] each imply the following.
@@
To make the connection to regular, normal Cartan geometries, we have to generalize the CR condition slightly to partially integrable, almost CR. A pseudoconvex, partially integrable, almost CR manifold is equivalent to a choice of regular, normal Cartan geometry modeled on the CR sphere $S^{2n+1}$, a homogeneous space for the action of $\mathrm{SU}(n+1,1)$ by CR diffeomorphisms. In particular, the group $\mathrm{SU}(n+1,1)$ has real rank one and its action on the sphere has a parabolic stabilizer subgroup. Frances [7] generalized the Ferrand-Obata and Schoen-Webster theorems, proving a Lichnerowicz-type result that applies to all real rank one parabolic geometries. Alt [1] modified Frances’s result to prove the following, under an appropriate Cartan-geometric definition of essentiality.
@@
There are four series of real rank one regular, normal parabolic geometries, corresponding to (1) conformal, (2) pseudoconvex, partially integrable, almost CR, (3) quaternionic contact, and (4) octonionic contact structures. These geometric structures are modeled on the homogeneous spaces (1) $\partial\mathbf{H}^{n+1}_{\mathbb{R}},$ (2) $\partial\mathbf{H}^{n+1}_{\mathbb{C}}$, (3) $\partial\mathbf{H}^{n+1}_{\mathbb{H}}$, and (4) $\partial\mathbf{H}^{2}_{\mathbb{O}}$, the boundaries of real, complex, quaternionic, and octonionic hyperbolic spaces.
@@
Lorentzian conformal manifolds of dimension $n\geq 3$ are regular, normal, parabolic Cartan geometries modeled on a homogeneous space for the group $\mathrm{O}(n,2).$ Since $\mathrm{O}(n,2)$ for $n\geq 3$ is simple of real rank $2,$ Lorentzian Lichnerowicz falls into an intermediate case between Frances and Alt’s Theorem 1.3 for real rank $1$ - a domain where Lichnerowicz-type results hold - and our Theorem 1.4 for real rank at least $3$ - a domain where they fail.
While this paper does not address failure of Lichnerowicz-type conjectures in real rank $2$, we believe that the Erickson techniques can be used to construct noncompact, nonflat, essential geometries for many but not all of of the Yamaguchi nonrigid parabolic model geometries modeled on homogeneous spaces for simple groups of real rank $2$. However, the construction of an essential compact quotient seems to rely in an important way on the assumption of real rank at least $3$.
@@ Acknowledgments
2 Lie Theory Background
2.1 Conventions
Given a real vector space $V$, we denote its complexification by $V^{\mathbb{C}}$. If $V$ is a complex vector space, $\overline{V}$ is the same vector space but with the conjugate complex structure. We call a real Lie algebra noncomplex it is not the underlying real Lie algebra of some complex Lie algebra. If $\mathfrak{g}$ is a Lie algebra and $\mathfrak{h}$ is a subalgebra, then $\mathfrak{z}(\mathfrak{g})$ represents the center of $\mathfrak{g}$ while $\mathfrak{z}_{\mathfrak{g}}(\mathfrak{h})$ represents the centralizer of $\mathfrak{h}$ in $\mathfrak{g}$. Given an inner product, a vector $v$ determines a covector $v_{\flat}$ and a covector $\alpha$ determines a vector $\alpha^{\sharp}$ via the musical isomorphisms. If $\alpha$ is a root or restricted root of a Lie algebra $\mathfrak{g}$, $\eta_{\alpha}$ refers to some nonzero vector in the rootspace or restricted rootspace $\mathfrak{g}_{\alpha}.$ The notation $\alpha_{k}$ refers to a simple root, while $\beta_{k}$ refers to a simple restricted root. For simple roots $\alpha_{i},\alpha_{j},$ the notation $(ij)$ refers to the composition of simple root reflections $s_{i}\circ s_{j}.$ The letter $\mu$ always refers to the highest root of a Lie algebra. Coefficients of the Cartan matrix are written $c_{ij}.$
2.2 Structure Theory of Real Lie Algebras
@@
$\mathfrak{g}=\mathfrak{z}_{\mathfrak{g}}(\mathfrak{a})\oplus\bigoplus_{\alpha\in\hat{\Delta}}\mathfrak{g}_{\alpha}$
@@
Given a simple system $\Delta^{0}$ for $\Delta^{+}$, the restrictions to $\mathfrak{a}$ of the roots in $\Delta^{0}\backslash\Delta_{c}$ form a simple system $\hat{\Delta}^{0}$ for $\hat{\Delta}^{+},$ with some pairs of simple roots in $\Delta^{0}$ restricting to a single simple restricted root in $\hat{\Delta}^{0}.$ Compact simple roots are excluded because they all restrict to $0.$ A subset of simple restricted roots $\hat{I}\subset\hat{\Delta}^{0}$ determines a parabolic subalgebra by the following process. Given a restricted root
@@
$\mathfrak{p}=\mathfrak{z}_{\mathfrak{g}}(\mathfrak{a})\oplus\bigoplus_{\alpha\in\hat{\Delta}:h_{\hat{I}}(\alpha)\geq 0}\mathfrak{g}_{\alpha}.$
Notably, $\mathfrak{p}=\mathfrak{g}$ if and only if $\hat{I}$$\varnothing$. Up to inner automorphism, this gives every parabolic subalgebra of $\mathfrak{g}.$ Therefore we will assume without loss of generality that a given parabolic subalgebra is in this form. Define
@@
$\mathfrak{g}_{0}:=\mathfrak{z}_{\mathfrak{g}}(\mathfrak{a})\oplus\bigoplus_{\alpha\in\hat{\Delta}(\mathfrak{g}_{0})}\mathfrak{g}_{\alpha},$
which is isomorphic to the quotient $\mathfrak{p}/\mathfrak{p}_{+}$ of $\mathfrak{p}$ by its nilradical. Since $\mathfrak{g}_{0}$ is reductive, its derived subalgebra $\mathfrak{g}_{0}^{ss}:=[\mathfrak{g}_{0},\mathfrak{g}_{0}]$ is semisimple. One possible source of confusion with the notation $\mathfrak{g}_{0}$ is that it can be mistaken for the restricted root space $\mathfrak{g}_{\alpha}$ when $\alpha=0.$ However, 0 is not considered a restricted root and this space is instead written $\mathfrak{z}_{\mathfrak{g}}(\mathfrak{a}).$
@@
It follows from Proposition 2.1 that $\beta_{k}^{\sharp}\in\mathfrak{g}_{0}^{ss}\cap\mathfrak{a}.$ From their definitions, we know that $\mathfrak{p}=\mathfrak{g}_{0}\oplus\mathfrak{p}_{+}.$ In fact, $h_{\hat{I}}$ induces a grading
$\mathfrak{g}=\mathfrak{g}_{-k}\oplus\cdots\oplus\mathfrak{g}_{0}\oplus\cdots\oplus\mathfrak{g}_{k},$
and we define $\mathfrak{g}_{-}:=\mathfrak{g}_{-k}\oplus\cdots\oplus\mathfrak{g}_{-1}$ and we already have $\mathfrak{p}_{+}=\mathfrak{g}_{1}\oplus\cdots\oplus\mathfrak{g}_{k},$ so that $\mathfrak{g}=\mathfrak{g}_{-}\oplus\mathfrak{g}_{0}\oplus\mathfrak{p}_{+}.$ There exists a special element $E\in\mathfrak{a}$ such that $[E,X]=iX$ for $X\in\mathfrak{g}_{i},$ known as the grading element. In particular, this definition implies $E\in\mathfrak{z}(\mathfrak{g}_{0}).$
@@
for some integers $n_{\beta}$. Then $\alpha|_{\mathfrak{a}}=\overline{\alpha}|_{\mathfrak{a}}.$ To draw the Satake diagram for a Lie algebra $\mathfrak{g},$ take the Dynkin diagram for $\mathfrak{g}^{\mathbb{C}},$ color the elements of $\Delta_{c}\cap\Delta^{0}$ black, and for noncompact simple roots $\alpha,$ if $\alpha\neq\overline{\alpha},$ connect $\alpha$ and $\overline{\alpha}$ by a bi-directional arrow. The resulting diagram is independent of a choice of maximally noncompact Cartan subalgebra $\mathfrak{c}$ and positive system $\Delta^{+}\subset\Delta$ compatible with $\sigma$. A full list of Satake diagrams of real Lie algebras $\mathfrak{g}$ such that $\mathfrak{g}^{\mathbb{C}}$ is simple is given in Appendix A.
@@ 2.3 Real and Complex Representations
If $\mathfrak{g}$ is complex semisimple and $X_{1},\dots X_{n}$ is a $\mathbb{C}$-basis for $\mathfrak{g},$ then $X_{1},\dots X_{n},iX_{1},\dots iX_{n}$ is an $\mathbb{R}$-basis for $\mathfrak{g}.$ Then given a $\mathbb{C}$-linear endomorphism $T:\mathfrak{g}\to\mathfrak{g},$ we have $\mathrm{tr}_{\mathbb{R}}(T)=2\mathrm{Re}(\mathrm{tr}_{\mathbb{C}}(T)).$ If $\mathfrak{g}$ is a complex Lie algebra, let $\langle\cdot,\cdot\rangle_{\mathbb{C}}$ be the Killing bracket on $\mathfrak{g}$, and let $\langle\cdot,\cdot\rangle_{\mathbb{R}}$ be the Killing bracket on $\mathfrak{g}$ as a real Lie algebra. Then $\langle x,y\rangle_{\mathbb{R}}=2\mathrm{Re}(\langle x,y\rangle_{\mathbb{C}}).$ Let $\mathfrak{h}_{0}\leq\mathfrak{h}$ be the subspace on which roots of $\mathfrak{g}$ are real-valued. Because $\langle\cdot,\cdot\rangle_{\mathbb{C}}$ is real-valued on $\mathfrak{h}_{0}$, if $x,y\in\mathfrak{h}_{0},$ then $\langle x,y\rangle_{\mathbb{R}}=2\langle x,y\rangle_{\mathbb{C}}.$ These brackets induce isomorphisms $\flat_{\mathbb{C}}:\mathfrak{h}_{0}\mapsto\mathfrak{h}_{0}^{*}$ and $\flat_{\mathbb{R}}:\mathfrak{h}_{0}\to\mathfrak{h}_{0}^{*},$ which both induce Killing brackets on $\mathfrak{h}_{0}^{*}.$ We have $\flat_{\mathbb{R}}=2\flat_{\mathbb{C}}.$ Calling their inverses $\sharp_{\mathbb{R}}$ and $\sharp_{\mathbb{C}},$ $\sharp_{\mathbb{C}}=2\sharp_{\mathbb{R}}.$
@@
Let $,\langle\cdot,\cdot\rangle_{\mathbb{C}}$$\langle\cdot,\cdot\rangle_{\mathbb{R}}$ be the complex and real Killing brackets induced on $\mathfrak{h}_{0}^{*}$. Then $\langle x,y\rangle_{\mathbb{C}}=2\langle x,y\rangle_{\mathbb{R}}.$
@@
Given $\mathfrak{g}$ real semisimple, a representation $\mathfrak{g}\curvearrowright V$, and $\alpha\in\mathfrak{a}^{*},$ a vector $v\in V$ is called a weight vector of weight $\alpha$ if $H\cdot v=\alpha(H)v$ for all $H\in\mathfrak{a}.$ The weight space $V_{\alpha}$ is the set of such vectors. A weight vector $v$ is called a lowest weight vector if it is annihilated by all negative restricted rootspaces of $\mathfrak{g}.$
Example 2.3.
Suppose $\mathfrak{g}$ is the underlying real Lie algebra of a complex semisimple Lie algebra. Let $\mathfrak{h}\leq\mathfrak{g}$ be the Cartan subalgebra. As explained in [2], $\mathfrak{h}\leq\mathfrak{g}$ is maximally noncompact with noncompact part $\mathfrak{h}_{0}\leq\mathfrak{h}$, the subspace on which all roots of $\mathfrak{g}$ are real valued. Then the restricted roots of $\mathfrak{g}$ are exactly the roots of $\mathfrak{g}$ restricted to $\mathfrak{h}_{0},$ and each restricted root space $\mathfrak{g}_{\alpha}$ for $\alpha\in\hat{\Delta}$ is a $2$ dimensional real space. Furthermore, if $\mathfrak{g}\curvearrowright V$ is a complex representation, the restricted weights of $\mathfrak{g}\curvearrowright V$ are the weights of $\mathfrak{g}\curvearrowright V$ restricted to $\mathfrak{h}_{0}.$
Given $\mathfrak{g}$ real semisimple, define the fundamental weights $\lambda_{i}\in\mathfrak{h}^{*}$ of $\mathfrak{g}^{\mathbb{C}}$ by the property $\langle\lambda_{i},\alpha_{j}\rangle=\delta_{ij}\frac{|\alpha_{j}|^{2}}{2}$ for all simple roots $\alpha_{j}\in\Delta^{0}.$ An integral weight $\gamma\in\mathfrak{h}^{*}$ decomposes as $\gamma=\sum\gamma^{i}\lambda_{i}$ for integers $\gamma^{i}.$ Define restricted fundamental weights $\hat{\lambda}_{i}\in\mathfrak{a}^{*}$ by the property $\langle\hat{\lambda}_{i},\beta_{j}\rangle=\delta_{ij}\frac{|\beta_{j}|^{2}}{2}$ for all simple restricted roots $\beta_{j}\in\hat{\Delta}^{0}$. A restricted weight $\sigma$ decomposes as $\sigma=\sum\sigma^{i}\hat{\lambda}_{i}$ for integers $\sigma^{i}.$
3 Cartan Geometry Background